# Tracing and Reversing Rank-One Model Edits

### v1

@@ Appendix C Additional Results
Figure 9: Increase in row-wise cosine similarity of $W_{n}$ after editing. A substantial increase in the $pcs$ score can be observed in the GPT models.
@@
Figure 11: The average pairwise cosine similarity ($pcs$) of edited and unedited matrices from different layers.
@@
Table 6: Pair-wise cosine similarity ($pcs$) scores with different relations from CounterFact.
C.1 Yago Dataset
@@
We use the knowledge base YAGO 4.5 Suchanek et al. (2024) to create an editing dataset that contains more diverse relations than CounterFact. YAGO 4.5 merges the taxonomy of Wikidata with the taxonomy of Schema.org to create a consistent knowledge base. We manually selected a set of diverse relations from YAGO and extracted the corresponding subject and object pairs. We filtered out relations with less than 1000 subject/object pairs. Afterwards we used DeepSeek R1 DeepSeek-AI et al. (2025) to generate editing prompts and paraphrased prompts for each relation. We manually checked the correctness of the generated prompts. YAGO 4.5 is licensed under a Creative Commons Attribution 4.0 International License (CC BY 4.0), and we will publish the dataset under the same license. We use 15 relations (shown in Tab. 4) from the newly constructed dataset in our experiments.
@@
Fig. 13 shows the direction distribution of row vectors in the update matrix. Fig. 14 shows the average pairwise cosine similarity ($pcs$) of edited and unedited matrices from different layers. Fig. 15 shows the increase in row-wise cosine similarity after editing. Tab. 7 shows the results for predicting the edited relation. Fig. 12 shows the accuracy of inferring the edited objects based on the edited weights.
@@
0 0.00 $\pm$ 0.00 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 95.71 $\pm$ 20.40 1.43 $\pm$ 11.95 91.43 $\pm$ 28.20
1 92.86 $\pm$ 25.94 5.71 $\pm$ 23.38 24.29 $\pm$ 43.19 61.43 $\pm$ 49.03 5.71 $\pm$ 23.38 82.86 $\pm$ 37.96
2 95.71 $\pm$ 20.40 2.86 $\pm$ 16.78 68.57 $\pm$ 46.76 10.00 $\pm$ 30.22 28.57 $\pm$ 45.50 41.43 $\pm$ 49.62
3 94.29 $\pm$ 23.38 1.43 $\pm$ 11.95 74.29 $\pm$ 44.02 8.57 $\pm$ 28.20 37.14 $\pm$ 48.67 34.29 $\pm$ 47.81
4 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 74.29 $\pm$ 44.02 8.57 $\pm$ 28.20 50.00 $\pm$ 50.36 20.00 $\pm$ 40.29
5 94.29 $\pm$ 23.38 0.00 $\pm$ 0.00 81.43 $\pm$ 39.17 4.29 $\pm$ 20.40 54.29 $\pm$ 50.18 17.14 $\pm$ 37.96
6 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 84.29 $\pm$ 36.66 4.29 $\pm$ 20.40 58.57 $\pm$ 49.62 17.14 $\pm$ 37.96
7 94.29 $\pm$ 23.38 0.00 $\pm$ 0.00 82.86 $\pm$ 37.96 0.00 $\pm$ 0.00 65.71 $\pm$ 47.81 10.00 $\pm$ 30.22
8 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 84.29 $\pm$ 36.66 0.00 $\pm$ 0.00 70.00 $\pm$ 46.16 8.57 $\pm$ 28.20
9 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 81.43 $\pm$ 39.17 0.00 $\pm$ 0.00 70.00 $\pm$ 46.16 7.14 $\pm$ 25.94
10 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 80.00 $\pm$ 40.29 0.00 $\pm$ 0.00 72.86 $\pm$ 44.79 7.14 $\pm$ 25.94
11 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 80.00 $\pm$ 40.29 0.00 $\pm$ 0.00 72.86 $\pm$ 44.79 7.14 $\pm$ 25.94
12 94.29 $\pm$ 23.38 0.00 $\pm$ 0.00 78.57 $\pm$ 41.33 0.00 $\pm$ 0.00 70.00 $\pm$ 46.16 8.57 $\pm$ 28.20
13 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 78.57 $\pm$ 41.33 0.00 $\pm$ 0.00 72.86 $\pm$ 44.79 7.14 $\pm$ 25.94
14 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 78.57 $\pm$ 41.33 1.43 $\pm$ 11.95 70.00 $\pm$ 46.16 7.14 $\pm$ 25.94
15 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 81.43 $\pm$ 39.17 1.43 $\pm$ 11.95 71.43 $\pm$ 45.50 5.71 $\pm$ 23.38
Table 8: Reversal/Editing accuracy on Yago with different $r-k$ approximations of $W^{\prime}_{V}$. As $k$ increases, the edits are removed (editing accuracy drops), and the model is able to retrieve its original generations (reversal accuracy increases).
@@
Table 9: Model outputs when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ on a random set of facts. We use the best $k$ for each model. The examples show that the model outputs with approximations (Approx. Output) are semantically close to the original/unedited outputs (Orig. Output).
@@
Figure 14: The average pairwise cosine similarity ($pcs$) of edited and unedited matrices from different layers (Yago Dataset).
Figure 15: Increase in row-wise cosine similarity of $W_{n}$ after editing with the Yago dataset. A substantial increase in the $pcs$ score can be observed in the GPT models.
Appendix D Implementation Details
Tab. 10 shows the dimensionality of the edited matrices in each model. Sec. D.1 describes the compute resources.
@@ Appendix D Implementation Details
CounterFact Meng et al. (2022) MIT License
YAGO Suchanek et al. (2024) CC BY 4.0
@@
D.1 Compute Resources
We ran all of our experiments were on an NVIDIA A100 GPU with 80GB. We started our experiments by editing the 3 models and storing the weight matrices. Editing took roughly 7 GPU days. The experiments for predicting the edits took roughly 2 GPU days. The experiments for inferring the edits from edited weights took roughly 7 GPU days. The experiments for reversing edits took roughly 5 GPU days.

### v2

@@ Appendix H Additional Results
Figure 15: Increase in row-wise cosine similarity of $W_{n}$ after editing. A substantial increase in the pcspcs score can be observed in the GPT models.
@@
Figure 17: The average pairwise cosine similarity (pcspcs) of edited and unedited matrices from different layers.
@@
Table 17: Pair-wise cosine similarity (pcspcs) scores with different relations from CounterFact.
H.1 Yago Dataset
@@
We use the knowledge base YAGO 4.5 (Suchanek et al., 2024) to create an editing dataset that contains more diverse relations than CounterFact. YAGO 4.5 merges the taxonomy of Wikidata with the taxonomy of Schema.org to create a consistent knowledge base. We manually selected a set of diverse relations from YAGO and extracted the corresponding subject and object pairs. We filtered out relations with less than 1000 subject/object pairs. Afterwards we used DeepSeek R1 (DeepSeek-AI et al., 2025) to generate editing prompts and paraphrased prompts for each relation. We manually checked the correctness of the generated prompts. YAGO 4.5 is licensed under a Creative Commons Attribution 4.0 International License (CC BY 4.0), and we will publish the dataset under the same license. We use 15 relations (shown in Tab. 15) from the newly constructed dataset in our experiments.
@@
Fig. 18 shows the direction distribution of row vectors in the update matrix. Fig. 19 shows the average pairwise cosine similarity (pcspcs) of edited and unedited matrices from different layers. Fig. 20 shows the increase in row-wise cosine similarity after editing. Tab. 18 shows the results for predicting the edited relation.
@@
0 0.00 97.14 0.00 95.71 1.43 91.43
1 92.86 5.71 24.29 61.43 5.71 82.86
2 95.71 2.86 68.57 10.00 28.57 41.43
3 94.29 1.43 74.29 8.57 37.14 34.29
4 95.71 0.00 74.29 8.57 50.00 20.00
5 94.29 0.00 81.43 4.29 54.29 17.14
6 97.14 0.00 84.29 4.29 58.57 17.14
7 94.29 0.00 82.86 0.00 65.71 10.00
8 95.71 0.00 84.29 0.00 70.00 8.57
9 95.71 0.00 81.43 0.00 70.00 7.14
10 97.14 0.00 80.00 0.00 72.86 7.14
11 97.14 0.00 80.00 0.00 72.86 7.14
12 94.29 0.00 78.57 0.00 70.00 8.57
13 95.71 0.00 78.57 0.00 72.86 7.14
14 97.14 0.00 78.57 1.43 70.00 7.14
15 95.71 0.00 81.43 1.43 71.43 5.71
Table 19: Reversal/Editing accuracy on Yago and ROME with different rkr-k approximations of $W^{\prime}_{V}$. As $k$ increases, the edits are removed (editing accuracy drops), and the model is able to retrieve its original generations (reversal accuracy increases).
@@
Table 20: Model outputs when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ on a random set of facts from the Yago dataset, edited with ROME. We use the best $k$ for each model. The examples show that the model outputs with approximations (Approx. Output) are semantically close to the original/unedited outputs (Orig. Output).
@@
Figure 19: The average pairwise cosine similarity (pcspcs) of edited and unedited matrices from different layers (Yago Dataset).
Figure 20: Increase in row-wise cosine similarity of $W_{n}$ after editing with the Yago dataset. A substantial increase in the pcspcs score can be observed in the GPT models.
@@
CounterFact (Meng et al., 2022) MIT License
YAGO (Suchanek et al., 2024) $C$ BY 4.0
@@
Appendix I The Use of LLMs
We used LLMs for polishing the writing, and for writing some parts of the code. We reviewed and validated all outputs. Additionally, LLMs were used to generate prompts for the Yago dataset as described in Sec. 3.